# Interpretable Visualizations of Data Spaces for Classification Problems

### v2

@@ 1 Introduction
Many approach QSPR through human-interpretable visualization via unsupervised machine-learning methods. Principal Component Analysis (PCA) is the most standard method for mapping tasks [5]. PCA involves identifying a linear (or non-linear, in the case of kernel PCA) combination of the features that maximizes the variance (i.e., information) retained. Beyond PCA, there exist a slue of other mapping techniques commonly used to qualitatively understand data spaces – t-distributed Stochastic Neighbor Embedding (t-SNE) [6], Uniform Manifold Approximation and Projection (U-MAP) [7], and Isometric Mapping (ISOMAP) [8] focus on the preservation of the local, global, and geometric data structure, respectively, and Locally Linear Embedding (LLE) [9] and multi-dimensional scaling (MDS) [10, 11] preserve distances between neighboring data points. While all of these methods excel at creating informative visualizations and identifying patterns in the structure of the data, the “P” (property) of “QSPR” is only spuriously accounted for, where any relationship between the visualized dimensions and the property of interest is purely coincidental.
@@
Thus, while methods from unsupervised learning and supervised learning address aspects of the QSPR problem, none can fully do so on their own. To address this problem, there exists a family of methods, known as hybrid learning, that draws elements from both supervised and unsupervised learning. Some of these methods include principal components regression [16], partial least squares regression [17], cluster-wise regression [18], continuum regression [19], and principal covariates regression (PCovR) [20]. Of these methods, PCovR is particularly valuable for QSPR because it is directly formulated as a multi-objective problem in which the regression and dimensionality reduction tasks are optimized simultaneously. PCovR has broad applications within the fields of climate science [21], economic forecasting [22], psychology [23], bioinformatics [24, 25, 26], and molecular crystallization [27]. Similar to PCA, PCovR can also be used to incorporate non-linear relationships through similarity kernels [28], and can be used with regularization to create sparse solutions [21, 25]. In addition, PCovR has been used to modify other techniques, such as clusterwise regression [29], farthest point sampling, and CUR decomposition [30].
@@ 2 Theory
2.1 Notation
We assume that the data is organized into two objects: a “feature matrix”, $\mathbf{X}$, that contains a set of input variables for a set of observations, and a “target matrix”, $\mathbf{Y}$, that contains the classes for each observation. The rows of the feature matrix $\mathbf{X}$ correspond to the individual samples and the columns correspond to the features, leading to shape $n_{\text{samples}}\times n_{\text{features}}$. The target matrix $\mathbf{Y}$ has shape $n_{\text{{samples}}}\times n_{\text{{labels}}}$ and contains the output labels for each observation, each of which can take an integer value between 00 and $n_{\text{classes}}-1$. We denote the latent-space projection of the input data as $\mathbf{T}$, which has shape $n_{\text{samples}}\times n_{\text{components}}$. Finally, we denote $\mathbf{Z}$ as the evidence tensor, which contains confidence scores for each class. $\mathbf{Z}$ has shape of $n_{\text{samples}}\times n_{\text{classes}}\times n_{\text{labels}}$. We assume that $\mathbf{P_{XZ}}$ is determined by a trained classification model, where, depending on the architecture of this model, $\mathbf{Z}$ is used to determine the class of the data points. In this context, we wish to learn linear mappings between $\mathbf{X}$, $\mathbf{T}$, and $\mathbf{Z}$, which we can define using projection matrices. In this work, we define $\mathbf{P}_{AB}$ as the projector from space $\mathbf{A}$ to space $\mathbf{B}$. Therefore, we can write $\mathbf{T}\equiv\mathbf{XP}_{XT}$ and $\mathbf{Z}=\mathbf{TP}_{TZ}=\mathbf{XP}_{XT}\mathbf{P}_{TZ}.$ The relationship between these objects is shown in Fig. 1.
@@ 2.2 Linear Classification
2.3 Principal Covariates Regression
@@
$\ell=\alpha\underbrace{||\mathbf{X-XP}_{XT}\mathbf{P}_{XT}^{T}||^{2}}_{\text{loss in reconstructing X}}+(1-\alpha)\underbrace{||\mathbf{Y}-\mathbf{XP}_{XT}\mathbf{P}_{TY}||^{2}}_{\text{loss in predicting Y}},$ (1)
@@
2.4 Principal Covariates Classification
@@ 3 Results and Discussion
3.1 What makes a molecule toxic? Exposing decision boundaries in toxicity classifications
@@
First, we compare the mapping results of PCovC as a function of our mixing parameter, $\alpha$, in Fig. 3. As expected, a PCA map of this dataset reflects little distinction between toxic and non-toxic molecules, where a subsequent analysis via logistic regression labels all data points as “non-toxic,” i.e., the majority population of the dataset. As $\alpha$ is lowered and the classification task bears more weight in Eq. (1), the activity cliff becomes clearer and our predictive power improves, even superceding that of logistic regression on the full-dimensional data. It is worth noting that the best performance, unlike PCovR, comes from a lower $\alpha$ value than 0.5 – this is likely due to the class imbalance within this dataset, and demonstrates the benefits of the tunability of PCovC.
@@
3.2 Parsing multi-class datasets: spectral analysis of organosulfur bonding environments
@@
Figure 6: Analysis of organosulfur dataset for different classification targets. Rows correspond to classification based on oxidation state (top), bond type (center), and aromaticity (bottom). (left) Underlying XANES spectra, colored by the class delineations. Heatmaps underneath the spectra correspond to correlations with the first two principal covariates. The main figure corresponds to the underlying spectra. (right) Maps according to PCA, LDA, and PCovC, colored by the class delineations. The PCA map will be the same for all three analyses. In the upper right, we indicate the accuracy of a logistic regression model on the resulting 2D map for the offset testing data. For PCovC, $\alpha$ is chosen to optimize test classification accuracy, indicated in each figure.
As can be seen in the first row of Fig. 6, it is straightforward to identify the the signals corresponding to different oxidation states, as the data separates well based on this set of classes. Because of this, a logistic regression model trained on the 2D PCA latent space yields perfect test accuracy, and LDA only misclassifies one test compound. Because PCA is well-suited to separate oxidation states, we can prioritize information retention in our PCovC mapping strategy. Thus, a larger $\alpha$ value of 0.85 is ideal to show the diversity of oxidation states. Note that this yields a map that is qualitatively similar to the PCA projection, with some minute differences.
Bond types are not well-separated in the PCA projection, given that the -2 oxidation state contains three different types. As shown in Fig. 6, the sulfides and the thiols are the most difficult to separate, leading to a loss in accuracy for both PCA (87%) and LDA (85%) . However, a logistic regression model trained on the PCovC latent space is still able to achieve a 92% bond type test classification accuracy. This demonstrates how incorporating classification loss into our mapping procedures “refocuses” our model; in purely unsupervised schema, having a wealth of data within one class versus another leads to this data dominating the mapping task. Because our sulfoxides and sulfones are easily recognizable, even a small amount of classification loss included in our mapping procedure (with $\alpha=0.95$) leads to significantly better separation of the thiocarbonyls, sulfides, and thiols.
The most difficult task is separating the compounds by aromaticity, as this is a property that cross-cuts the “higher-order” classes. As is suggested by the maps in Fig. 6, the principal components seem to be capturing the oxidation state and bond type of the organosulfur compounds. However, because neither of these properties has a direct relationship with bond aromaticity, a logistic regression model trained from the PCA latent space struggles (accuracy of 48%). However, PCovC and LDA are each able to find latent spaces in which the compounds are well delineated by their aromaticity, with accuracies of 78% and 73%, respectively.
@@
3.3 Understanding feature importance in non-linear classification of metals/nonmetals
@@
Figure 7: Projections of inorganic materials found in [41] using Kernel PCA, KDA, and Kernel PCovC. Color and opacity denote metal (blue) versus non-metal (red) and training (transparent) versus testing (opaque) splits, respectively. Above each map is the distribution of points in each class.
@@
Figure 8: Correlations of materials characteristics with the non-linear KPCovC and KPCA maps. (left) Absolute value of Pearson correlation coefficients for different input features with the resulting maps from Fig. 7. Darker squares denote stronger correlation. (right) Recoloring of KPCovC (main) and KPCA (inset) by feature values, from low (red) to high (blue) on a linear scale.
@@
4 Conclusions
@@ A.1 Principal Covariates Classification
A.1.1 Inner Product of Evidence Tensors

### v3

@@ I Introduction
Many approach QSPR through human-interpretable visualization via unsupervised machine-learning methods. Principal Component Analysis (PCA) is the most standard method for mapping tasks [5]. PCA involves identifying a linear (or non-linear, in the case of kernel PCA) combination of the features that maximizes the variance (i.e., information) retained. Beyond PCA, there exist a slue of other mapping techniques commonly used to qualitatively understand data spaces – t-distributed Stochastic Neighbor Embedding (t-SNE) [6], Uniform Manifold Approximation and Projection (UMAP) [7], and Isometric Mapping (ISOMAP) [8] focus on the preservation of the local, global, and geometric data structure, respectively, and Locally Linear Embedding (LLE) [9] and multi-dimensional scaling (MDS) [10, 11] preserve distances between neighboring data points. While all of these methods excel at creating informative visualizations and identifying patterns in the structure of the data, the “P” (property) of “QSPR” is only spuriously accounted for, where any relationship between the visualized dimensions and the property of interest is purely coincidental.
@@
Thus, while methods from unsupervised learning and supervised learning address aspects of QSPR, none can fully do so on their own. There has been an emergence of "visualization as interpretation" techniques that aim to provide a solution. Briefly, the vast majority of these methods aim to compute invertible mappings that represent the partitioning and clustering of data in higher-dimensional spaces [15, 16, 17, 18]. For example, DeepView, the most notable of these methods, combines UMAP and Fisher distance for projection and then samples the resultant latent space on a grid to assess data boundaries[19]. However, computational cost and model complexity have remained large challenges for many of these methods, particularly those reliant on multi-layer architectures or grid-based assessments. And, methods based on high-non-linear operations or architectures can result in topologically-bound latent spaces, where arbitrary points or pixels in a grid may not directly correspond to physically realizable data points [20].
We address these challenges by exploiting a family of methods, known as hybrid learning, that draws elements from both supervised and unsupervised learning. Some of these methods include principal components regression [21], partial least squares regression [22], cluster-wise regression [23], continuum regression [24], and principal covariates regression (PCovR) [25]. Of these methods, PCovR is particularly valuable for QSPR because it is directly formulated as a multi-objective problem in which the regression and dimensionality reduction tasks are optimized simultaneously. PCovR has broad applications within the fields of climate science [26], economic forecasting [27], psychology [28], bioinformatics [29, 30, 31], and molecular crystallization [32]. Similar to PCA, PCovR can also be used to incorporate non-linear relationships through similarity kernels [33], and can be used with regularization to create sparse solutions [26, 30]. In addition, PCovR has been used to modify other techniques, such as clusterwise regression [34], farthest point sampling, and CUR decomposition [35].
@@ II Theory
II.1 Notation
We assume that the data is organized into two objects: a “feature matrix”, $\mathbf{X}$, that contains a set of input variables for a set of observations, and a “target matrix”, $\mathbf{Y}$, that contains the classes for each observation. The rows of the feature matrix $\mathbf{X}$ correspond to the individual samples and the columns correspond to the features, leading to shape $n_{\text{samples}}\times n_{\text{features}}$. The target matrix $\mathbf{Y}$ has shape $n_{\text{{samples}}}\times n_{\text{{labels}}}$ and contains the output labels for each observation, each of which can take an integer value between 0 and $n_{\text{classes}}-1$. We denote the latent-space projection of the input data as $\mathbf{T}$, which has shape $n_{\text{samples}}\times n_{\text{components}}$. Finally, we denote $\mathbf{Z}$ as the evidence tensor, which contains confidence scores for each class. $\mathbf{Z}$ has shape of $n_{\text{samples}}\times n_{\text{classes}}\times n_{\text{labels}}$. We assume that $\mathbf{P_{XZ}}$ is determined by a trained classification model, where, depending on the architecture of this model, $\mathbf{Z}$ is used to determine the class of the data points. In this context, we wish to learn linear mappings between $\mathbf{X}$, $\mathbf{T}$, and $\mathbf{Z}$, which we can define using projection matrices. In this work, we define $\mathbf{P}_{AB}$ as the projector from space $\mathbf{A}$ to space $\mathbf{B}$. Therefore, we can write $\mathbf{T}\equiv\mathbf{XP}_{XT}$ and $\mathbf{Z}=\mathbf{TP}_{TZ}=\mathbf{XP}_{XT}\mathbf{P}_{TZ}.$ The relationship between these objects is shown in Fig. 1.
@@ II.2 Linear Classification
II.3 Principal Covariates Regression
@@
[…]
@@
II.4 Principal Covariates Classification
@@ III Results and Discussion
III.1 What makes a molecule toxic? Exposing decision boundaries in toxicity classifications
@@
First, we compare the mapping results of PCovC as a function of our mixing parameter, $\alpha$, in Fig. 3. As expected, a PCA map of this dataset reflects little distinction between toxic and non-toxic molecules, where a subsequent analysis via logistic regression labels all data points as “non-toxic,” i.e., the majority population of the dataset. As shown in the SI, using nonlinear unsupervised dimensionality reduction algorithms such as t-SNE or UMAP also fails to distinguish toxic molecules from non-toxic molecules. As $\alpha$ is lowered and the classification task bears more weight in Eq. (1), the activity cliff becomes clearer and our predictive power improves, even superceding that of logistic regression on the full-dimensional data. It is worth noting that the best performance, unlike PCovR, comes from a lower $\alpha$ value than 0.5 – this is likely due to the class imbalance within this dataset, and demonstrates the benefits of the tunability of PCovC.
@@
III.2 Parsing multi-class datasets: spectral analysis of organosulfur bonding environments
@@
Figure 6: Analysis of organosulfur dataset for different classification targets. Rows correspond to classification based on oxidation state (top), bond type (center), and aromaticity (bottom). (left) Underlying XANES spectra, colored by the class delineations. The main figure corresponds to the underlying spectra. (right) Maps according to PCA, LDA, and PCovC, colored by the class delineations. The PCA map will be the same for all three analyses. In the upper right, we indicate the accuracy of a logistic regression model on the resulting 2D map for the offset testing data. For PCovC, $\alpha$ is chosen to optimize test classification accuracy, indicated in each figure.
As can be seen in the first row of Fig. 6, it is straightforward to identify the the signals corresponding to different oxidation states, as the data separates well based on this set of classes. Because of this, a logistic regression model trained on the 2D PCA latent space yields perfect test accuracy, and LDA only misclassifies one test compound. Because PCA is well-suited to separate oxidation states, the choice of $\alpha$ for a PCovC map is somewhat arbitrary, hence we choose $\alpha=0.5$. Note that the map shown is qualitatively similar to the PCA projection, with some minute differences.
Bond types are not well-separated in the PCA projection, given that the -2 oxidation state contains three different types. As shown in Fig. 6, the sulfides and the thiols are the most difficult to separate, leading to a loss in accuracy for both PCA (87%) and LDA (85%) . However, a logistic regression model trained on the PCovC latent space is still able to achieve a 92% bond type test classification accuracy. This demonstrates how incorporating classification loss into our mapping procedures “refocuses” our model; in purely unsupervised schema, having a wealth of data within one class versus another leads to this data dominating the mapping task.
The most difficult task is separating the compounds by aromaticity, as this is a property that cross-cuts the “higher-order” classes. As is suggested by the maps in Fig. 6, the principal components seem to be capturing the oxidation state and bond type of the organosulfur compounds. However, because neither of these properties has a direct relationship with bond aromaticity, a logistic regression model trained from the PCA latent space struggles (accuracy of 48%). However, PCovC and LDA are each able to find latent spaces in which the compounds are well delineated by their aromaticity, with accuracies of 77% and 73%, respectively.
Figure 7: Projections of inorganic materials found in [46] using Kernel PCA, KDA, and Kernel PCovC, and the corresponding $\alpha$ analysis for this non-linear mapping. Color and opacity denote metal (blue) versus non-metal (red) and training (transparent) versus testing (opaque) splits, respectively. Above each map is the distribution of points in each class.
@@
III.3 Understanding feature importance in non-linear classification of metals/nonmetals
@@
Figure 8: Correlations of materials characteristics with the non-linear KPCovC and KPCA maps. (left) Absolute value of Pearson correlation coefficients for different input features with the resulting maps from Fig. 7. Darker squares denote stronger correlation. (middle) Ranked SHAP values and random forest feature importance. Darker squares indicate a higher importance. (right) Recoloring of KPCovC (main) and KPCA (inset) by feature values, from low (red) to high (blue) on a linear scale.
@@
In general, the SHAP values and random forest feature importances correlate strongly with the representation of the features in the first principal covariate. However, interpreting models using these techniques has limitations. First, for these feature importance metrics to be useful, it is necessary that the features themselves can be interpreted. While this is not the case here, when abstract descriptors are used, such as the RDKit fingerprint in it will be difficult to relate these metrics to chemical concepts, which will make QSPR unachievable. With PCovC, model interpretation does not have to be limited to features the model uses for training, as visualizing or computing correlation coefficients can be performed trivially with the PCovC latent space. In addition, it is difficult to determine the span of the feature’s importance across the space of observations using solely feature importance metrics. For example, consider the map showing the average number of p electrons in the KPCovC projection in 8. This map shows that there is a strong population of metals with a small number of p electrons, but this population does not span the entire population of metals. While SHAP can similarly achieve sample-level resolution in this manner, PCovC can also be used to identify subgroups across the decision boundary, such as the sets of materials with high and low electronegativities in Fig. 8.
III.4 Enhancing nonlinear dimensionality reduction techniques with MNIST
This slew of methods, as demonstrated in [47, 48], extend in usefulness beyond those problems that are chemically-motivated. Furthermore, they offer benefits to traditional mapping workflows. To demonstrate both of these points, we conclude with a two-step workflow on the benchmark dataset MNIST.
Nonlinear methods are often essential when facing problems in which local correlations between features dictate class labels. Consequently, with the MNIST handwritten digits dataset, dimensionality reduction algorithms that optimize the preservation of local neighborhoods of points in their reduced spaces, such as t-SNE and UMAP, can find embeddings with far cleaner digit separation than PCA or even LDA [7]. This advantage, however, is balanced by an increased computational cost. For this reason, when working with large datasets, many have considered using PCA as a preprocessing step to reduce data and make nonlinear dimensionality reduction feasible [47, 48]. At the same time, linear classification models can achieve a respectable performance with MNIST; despite its simplicity, a logistic regression model trained on the flattened images can achieve an accuracy of 92%. This begs the question: if there exists linearly decodable information that can be used to partially separate classes, can this information be used to enhance a nonlinear mapping strategy?
Figure 9: PCovC and PCA-reduced UMAP projections on the MNIST dataset. (top) Accuracy of a logistic regression model trained in PCA, PCovC, PCovC-reduced UMAP, and PCA-reduced UMAP latent spaces. (middle row) UMAP projections of MNIST obtained from PCA-reduced latent spaces. (bottom row) UMAP projections of MNIST obtained from PCovC-reduced latent spaces. In the titles, we indicate the accuracy of a logistic regression model on the resulting 2D map for the testing data. Color corresponds to the digit label and the backgrounds show the estimated decision boundaries.
In Fig. 9, we explore the use of PCovC as a replacement to PCA as a preprocessing step to UMAP. We observe that UMAP projections learned from PCovC-reduced subspaces of the data result in clearer digit separation than those learned from PCA-reduced subspaces. In practice, this means that PCovC can be used in two ways: PCovC can substitute PCA at a fixed subspace dimensionality to enhance the quality of the subsequent UMAP embedding without a significant increase in the computational cost, or it can be used to achieve stronger dimensionality reduction prior to UMAP without sacrificing class delineation in the resulting projection.
IV Conclusions
@@ A.1 Principal Covariates Classification
A.1.1 Computational Scaling of PCovC
The primary cost of performing PCovC lies in the calculation and subsequent eigendecomposition of the modified Gram matrix $\mathbf{\tilde{K}}$ or modified Covariance matrix $\mathbf{\tilde{C}}$. For a feature matrix $\mathbf{X}$ of dimension $m\times n$, the explicit calculation of $\mathbf{\tilde{K}}$ or $\mathbf{\tilde{C}}$ will have complexity $\mathcal{O}(m^{2}n)$ or $\mathcal{O}(mn^{2})$, respectively. The equivalence of their eigendecompositions implies this can be done on the option corresponding to the lesser of $m$ and $n$. The scaling of the eigendecomposition is $\mathcal{O}(kmn)$, where $k$ is the number of eigenvectors we choose to retain. The overall complexity for PCovC will therefore be $\mathcal{O}(mn^{2}+kmn)$, where we note that the hidden prefactor for the eigendecomposition is typically much larger than that of matrix multiplication due to the latter’s parallelizability.
A.1.2 Inner Product of Evidence Tensors
@@ A.2.3 Metal/Nonmetal Classification
A.2.4 MNIST
The MNIST dataset [52] was obtained via the tensorflow API. MNIST contains a training set of 70,000 handwritten digits (60,000 train, 10,000 test) in the form of 28x28 grayscale images. Before applying dimensionality reduction, we first flatten each of the images to form a feature vector of length 784, then scale the intensity of each pixel in the image to have zero mean and unit variance.